OpenStudy (anonymous):

Two similar cylinders have heights 6 inches and 10 inches. What is the ratio of their surface areas? 3 : 5 9 : 25 1 : 4 9 : 10

OpenStudy (mathstudent55):

The total surface area of a cylinder is the lateral surface plus the area of the two bases. A = 2(pi)rh + 2(pi)r^2 In you case for the one with height 6: A6 = 2(pi)r(6) + 2(pi)r^2 For the one with height 10 A10 = 2(pi)r(10) + 2(pi)r^2 The ratio (sfter simplifying is A6/A10 = [12(pi)r + 2(pi)r^2] / [20(pi)r + 2(pi)r^2] Now you can factor out 2, r, and pi from both the numerator and denominator A6/A10 = [2(pi)r (6 + r)] / [2(pi)r(10 + r)] The 2(pi)r cancels out, and you're left with A6/A10 = (r + 6)/(r + 10)

OpenStudy (anonymous):

what? so how do i get the answer?

OpenStudy (mathstudent55):

I see this is not a choice. That means your problem wants only the lateral area without the bases. I'll do it next.

OpenStudy (anonymous):

would the answer be 3:5? cuz half of 6 and 10 are 2 and 5

OpenStudy (mathstudent55):

Lateral area is 2(pi)rh A6 = 2(pi)r(6) = 12(pi)r A10 = 2(pi)r(10) = 20(pi)r A6/A10 = [12(pi)r]/{20(pi)r] pi and r cancels out A6/A10 = 12/20 which reduces to 3/5

OpenStudy (mathstudent55):

You are correct, 3:5

OpenStudy (anonymous):

ok thank you! :)